### v1

# Optimal discrimination of quantum sequences

Abstract
A key concept of quantum information theory is that accessing information encoded in a quantum system requires us to discriminate between several possible states the system could be in. A natural generalization of this problem, namely, quantum sequence discrimination, appears in various quantum information processing tasks, the objective being to determine the state of a finite sequence of quantum states. Since such a sequence is a composite quantum system, the fundamental question is whether an optimal measurement is local, i.e., comprising measurements on the individual members, or collective, i.e. requiring joint measurement(s). In some known instances of this problem, the optimal measurement is local, whereas in others, it is collective. But, so far, a definite prescription based solely on the problem description has been lacking. In this paper, we prove that if the members of a given sequence are drawn secretly and independently from an ensemble or even from different ensembles, the optimum success probability is achievable by fixed local measurements on the individual members of the sequence, and no collective measurement is necessary. This holds for both minimum-error and unambiguous state discrimination paradigms.
One of the characteristic features of quantum theory is that composite quantum systems can possess nonlocal properties. This is often associated with entangled systems violating a Bell inequality bell1966problem ; brunner2014bell . However, an unentangled system, whose parts may not have interacted in the past, can also exhibit nonlocality, conceptually different from the Bell type peres1991optimal ; bennett1999quantum ; wootters2006distinguishing ; chitambar2013revisiting ; halder2019strong . Specifically, when determining the state of an unentangled system, known to be in one of several possible states, a joint measurement of the whole system is sometimes necessary. Thus a fundamental question in quantum information is when does optimal extraction of information from a composite system demands access to the whole and when it does not.
This question frequently arises in many quantum information theoretic protocols of practical importance, such as quantum change point detection sentis2016quantum ; sentis2017exact ; mohan2023generalized , multiple copy state discrimination higgins2011multiple , and quantum key distribution cerutti2024optimal . In all these scenarios, essentially, the objective is to determine the state of an unknown quantum sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ whose members $\rho_{i}$ are drawn from ensembles of which we have complete knowledge. For a given sequence of finite length, this boils down to the problem of sequence discrimination, a natural generalization of the well-studied quantum state discrimination problem bae2015quantum ; barnett2009quantum ; chefles2000quantum ; bergou200411 , where one aims to discriminate between the possible states a quantum system could be in. Since the state of a sequence $\left(\rho_{1},\dots,\rho_{k}\right)$ is of the form $\rho_{1}\otimes\cdots\otimes\rho_{k}$, the question is whether the optimal measurement, optimal according to some well-defined figure of merit, is local, i.e., comprising measurements on the individual members or collective, i.e. requiring joint measurement(s). Moreover, even when local access is sufficient, it is necessary to know whether coordinated or adaptive strategies provide any advantage over fixed strategies that do not involve adaptation based on the outcomes of already performed measurements. In fact, we know of instances where the optimal measurement is local and fixed gupta2024unambiguous , local and adaptive acin2005multiple , and collective calsamiglia2010local ; in particular, in the quantum change-point problem, it is collective sentis2016quantum , whereas in multiple-copy discrimination, depending on the problem specification, it could be either of the two acin2005multiple ; calsamiglia2010local .
Unfortunately, given a sequence discrimination problem, we do not have any definite characterization or prescription that could, at the very least, tell us about the nature of the optimal measurement. As it is, discriminating between states of a composite quantum system is known to be challenging wootters2006distinguishing ; chitambar2013revisiting ; halder2019strong and for sequences more so as they could be of any finite length and also come in different varieties: they could be i.i.d., where the members of a sequence are independent (i.e., selection of an individual member does not depend on what states have already appeared in the earlier positions) and identically distributed (i.e., the members are drawn from the same ensemble), or non-i.i.d., where members are either drawn independently but from different ensembles or not independently at all.
In this paper, we take a significant step towards addressing whether an optimal measurement is local or collective only from the problem description. We prove that if the members of a sequence are selected secretly and independently either from the same ensemble or even from different ensembles, the optimal measurement is local; in particular, it constitutes fixed measurements on the individual elements. The result holds for both minimum-error and unambiguous state discrimination paradigms. For the latter, however, additional assumptions are required for a nontrivial result chefles1998unambiguous ; feng2004unambiguous .
The main implication of our result is the following. We have shown that if the members of a given sequence are selected independently, irrespective of the probability distribution associated with the selection being identical or different, the optimal measurement to determine its state is always local and fixed. Therefore, if the independence condition does not hold, all types of measurements mentioned earlier, are possible. Future works, therefore, only need to address the optimality question in the non-independent scenario.
Suppose we are given an unknown sequence of length $k\in\mathbb{N}$ and wish to determine its state as well as possible. The members of the given sequence are secretly and independently drawn from potentially different but known ensembles. The ensemble from which the $i$-th member is drawn is labelled as $\mathcal{E}^{i}$, where
$\mathcal{E}^{i}=\left\{\left(\eta_{j}^{i},\rho_{j}^{i}\right):j=1,\dots,\ell_{i}\right\}$ (1)
with $\rho_{j}^{i}$ being density operators on $\mathbb{C}^{d_{i}}$ for $d_{i}\geqslant 2$. Thus, for a sequence of length $k$, we have $k$ such ensembles $\mathcal{E}^{1},\dots,\mathcal{E}^{k}$. This way, we can account for all possibilities: every member is drawn from a different ensemble; some are drawn from the same ensemble, in which case, though the labelling of ensembles is different, they are, in fact, identical; every member is drawn from the same ensemble, so here no indexing of ensembles would be necessary. Note that the last one corresponds to i.i.d. sequences, whereas every other case corresponds to independent but not identically distributed sequences.
Thus, given a sequence of length $k$, we only know that the $i$-th member is drawn from $\mathcal{E}^{i}$ and $\eta_{j}^{i}$ is the prior probability of it being $\rho_{j}^{i}$. With this, the ensemble of all possible sequences of length $k$ is given by
$\mathcal{E}_{k}=\left\{\left(\eta_{x_{1}}^{1}\cdots\eta_{x_{k}}^{k},\rho_{x_{1}}^{1}\otimes\cdots\otimes\rho_{x_{k}}^{k}\right)\right\},$ (2)
where each $x_{i}\in\left[\ell_{i}\right]$ for all $i\in\left[k\right]$. A simple counting argument shows that $\mathcal{E}_{k}$ contains $\ell=\prod_{i=1}^{k}\ell_{i}$ sequences. The objective is to determine whether the optimal measurement discriminating between the elements of $\mathcal{E}_{k}$ is collective or local.
The optimality of a measurement subject to a given set of states depends on the choice of measurement strategy. Here, we consider both minimum-error and unambiguous state discrimination paradigms. The former minimizes the average error and applies to any set of states. The corresponding measure is the success probability, the maximum probability that the unknown state is correctly determined holevo1973statistical ; helstrom1969quantum . The latter ivanovic1987differentiate ; dieks1988overlap ; peres1988differentiate , however, applies to sets of states that satisfy a specific condition, and if this condition is met, this approach correctly determines the unknown state with a nonzero probability; for example, a set of pure states can be unambiguously discriminated if and only if they are linearly independent chefles1998unambiguous . Closed-form solutions for both are known in the two-state case helstrom1969quantum ; jaeger1995optimal and in specific instances where the states satisfy certain symmetry properties sun2001optimum . For generic ensembles, finding solutions is generally hard; however, semidefinite programs exist eldar2003designing ; eldar2003semidefinite , and the optimum success probability can be obtained as the output of such programs.
In what follows, we first state the main result and then go on to prove it for both minimum-error and unambiguous discrimination.
Theorem 1.
Let $p\left(\mathcal{E}\right)$ denote the optimum probability for minimum-error or unambiguous discrimination between the elements of an ensemble $\mathcal{E}$. Then
$p\left(\mathcal{E}_{k}\right)=\prod_{i=1}^{k}p\left(\mathcal{E}^{i}\right).$
Thus the optimal measurement for discriminating between the elements of $\mathcal{E}_{k}$ is local and comprises of individual measurements on the members of the sequence. In particular, for the $i$-th member the local measurement corresponds to the optimal measurement that discriminates between the elements of the ensemble $\mathcal{E}^{i}$. Note that, for unambiguous discrimination, a nontrivial result, i.e., $p\left(\mathcal{E}_{k}\right)>0$ is obtained provided for each ensemble $\mathcal{E}^{i}$, the corresponding elements can be unambiguously discriminated.
Minimum-error discrimination of quantum sequences.—We need the following two results to prove Theorem 1 for this case. The first one is the Holevo–Yuen–Kennedy–Lax Theorem yuen1975optimum ; watrous2018theory , which provides a necessary and sufficient condition for the optimality of a minimum-error discrimination measurement for a given ensemble (note that the optimal measurement is not unique bae2013structure ). We only state the theorem here, the proof can be found in watrous2018theory .
Theorem 2 (Holevo–Yuen–Kennedy–Lax yuen1975optimum ; watrous2018theory ).
Let $N$ be a positive integer and $\mathcal{X}$ be a complex Euclidean space of finite dimension. Given an ensemble =$\mathcal{E}=\{(q_{i},\sigma_{i}):i\in[N]\}$ of density operators on $\mathcal{X}$, a measurement $\{M_{1},\ldots,M_{N}\}$ is optimal for minimum-error discrimination of the elements of $\mathcal{E}$ if and only if
$\sum_{i=1}^{N}q_{i}\sigma_{i}M_{i}\succeq q_{j}\sigma_{j}$ (3)
for all $j\in[N]$.
The second result we need is the following lemma, whose proof can be found in the supplemental material (SM).
Lemma 3.
Let $A,B,C,$ and $D$ be positive semidefinite operators such that $A\succeq C$ and $B\succeq D$. Then $A\otimes B\succeq C\otimes D$.
We now present the proof of Theorem 1 for minimum-error discrimination.
Proof.
Without loss of generality, assume that an optimal measurement for minimum-error discrimination between the elements of the ensemble $\mathcal{E}^{i}$ is $M^{i}=\{M_{x}^{i}\}_{x=1}^{\ell_{i}}\text{ where }M^{i}_{x}\in\operatorname{Pos}(\mathbb{C}^{d_{i}})\text{ and }\sum_{x=1}^{\ell_{i}}M_{x}^{i}=\mathbb{1}_{d_{i}}$ for all $x\in[\ell_{i}]$ and $i\in[k]$, where $\operatorname{Pos}(\mathbb{C}^{d_{i}})$ denotes the set of positive semidefinite operators on $\mathbb{C}^{d_{i}}$. It follows that
$\sum_{x=1}^{\ell_{i}}\eta^{i}_{x}\rho^{i}_{x}M^{i}_{x}\succeq\eta^{i}_{y}\rho^{i}_{y}$ (4)
for all $y\in[\ell_{i}]$ and $i\in[k]$. This represents a set of $\ell_{i}$ conditions, one for each state in $\mathcal{E}^{i}$. We select $k$ of these conditions, one for each $i\in[k]$. Our selection can be denoted by a $k$-tuple $(y_{1},\dots,y_{k})$, which means that we are considering inequality (4) for the states $\rho^{1}_{y_{1}},\dots,\rho^{k}_{y_{k}}$.
Taking the tensor product of the operators on both sides of these inequalities and applying Lemma 3 we get
$\bigotimes_{i=1}^{k}\sum_{x_{i}=1}^{\ell_{i}}\eta^{i}_{x_{i}}\rho^{i}_{x_{i}}M^{i}_{x_{i}}\succeq\bigotimes_{i=1}^{k}\eta^{i}_{y_{i}}\rho^{i}_{y_{i}},$ (5)
where $y_{i}\in[\ell_{i}]$ for all $i\in[k]$. The right side of the above inequality is the sequence $\rho^{1}_{y_{1}}\otimes\dots\otimes\rho^{k}_{y_{k}}$. By selecting all possible $k$-tuples from inequality (4) we will have a collection of conditions, one for each sequence, which is precisely what inequality (5) represents. The left-hand side of (5) can be expanded as
$\sum_{x_{1},\ldots,x_{k}}\eta^{1}_{x_{1}}\cdots\eta^{k}_{x_{k}}(\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}})(M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}).$
As $M^{i}=\{M_{x}^{i}\}_{x=1}^{\ell_{i}}$ forms a measurement on the ensemble $\mathcal{E}^{i}$, we observe that
$\sum_{x_{1},\ldots,x_{k}}M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}=$ $\sum_{x_{1}=1}^{\ell_{1}}M^{1}_{x_{1}}\otimes\cdots\otimes\sum_{x_{k}=1}^{\ell_{k}}M^{k}_{x_{k}}$ (6)
$=$ $\mathbb{1}_{d_{1}}\otimes\cdots\otimes\mathbb{1}_{d_{k}},$
showing that $\{M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}\}$ is a bonafide measurement on $\mathcal{E}_{k}$. This, together with equation (5), demonstrates that the measurement whose elements are ${M^{1}_{x_{1}}\otimes\cdots\otimes M^{k}_{x_{k}}}$ is indeed an optimal measurement for discriminating the elements of the ensemble $\mathcal{E}_{k}$, as established by Theorem 2.
Therefore, the optimal measurement for discriminating the sequences can be achieved by performing the optimal measurement for each ensemble on its corresponding state. As a result, the probability of correctly identifying the entire sequence is the product of the probabilities of correctly identifying each individual state within the sequence. ∎
Unambiguous discrimination of quantum sequences.— Unlike the minimum-error case, for non-trivial unambiguous discrimination, a set of quantum states $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ has to satisfy the following condition.
Let $\text{supp}(Q)$ denote the Hilbert space spanned by the eigenvectors of the matrices $\{\sigma_{1},\ldots,\sigma_{N}\}$ that correspond to nonzero eigenvalues. Additionally, let $S(\mathcal{E})$ represent the set of states of an ensemble $\mathcal{E}$.
Lemma 4.
(from feng2004unambiguous ) The set of density operators $Q=\{\sigma_{1},\ldots,\sigma_{N}\}$ can be unambiguously discriminated if and only if $\text{supp}(Q)\neq\text{supp}(Q_{i})$ where $Q_{i}=Q\setminus\{\sigma_{i}\}$ for all $i\in[N]$. If $\sigma_{i}$ are rank $1$ operators (pure states), then this condition is same as the set $Q$ being linearly independent.
Since we now want to discriminate the set of sequences unambiguously, they must satisfy Lemma 4. That is, for $S(\mathcal{E}_{k})$ to be unambiguously discriminable it must hold that
$\text{supp}(S(\mathcal{E}_{k}))\neq S(\mathcal{E}_{k})\setminus\{\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}\}$ (7)
for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. In the following lemma (proof in SM), we show that a set of quantum sequences satisfies this condition if and only if the individual ensembles do.
Lemma 5.
The set of quantum sequences satisfies the condition $\text{supp}(S(\mathcal{E}_{k}))\neq S(\mathcal{E}_{k})\setminus\{\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}\}$ for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$ if and only if the ensembles $\mathcal{E}^{i}$ satisfy $\text{supp}(S(\mathcal{E}^{i}))\neq\text{supp}(S(\mathcal{E}^{i}))\setminus\{\rho^{i}_{j}\}$ for all $i\in[k]$ and $j\in[\ell_{i}]$.
The following theorem is immediate from Lemmas 4 and 5.
Theorem 6.
The elements of $S(\mathcal{E}_{k})$ can be unambiguously discriminated if and only if the elements of $S(\mathcal{E}^{i})$ can be unambiguously discriminated for each $i$.
We now present the proof of Theorem 1 for the unambiguous case. The unambiguous discrimination problem of mixed quantum states can be cast as a semidefinite program (SDP) eldar2004optimal . Let us first present this SDP for an ensemble of states $\mathcal{E}=(Q,q)$, where $Q=\{\sigma_{i}\}_{i=1}^{N}$ is a set of density operators acting on $\mathbb{C}^{d}$ and $q=(q_{1},\ldots,q_{N})$ is the vector of a priori probabilities of states of $Q$. Without loss of generality, we can assume that the eigenvectors of the states of $Q$ that correspond to nonzero eigenvalues span $\mathbb{C}^{d}$. This is analogous to assuming that for unambiguous discrimination of pure states, we can take the dimension of the Hilbert space they live in to equal the dimension of the space spanned by them. We define $\widetilde{Q_{i}}$ as the intersection of all kernels $\mathcal{K}_{j}$ of $\sigma_{j}$, excluding $\mathcal{K}_{i}$. This can be expressed mathematically as $\widetilde{Q_{i}}=\cap_{j=1,j\neq i}^{N}\mathcal{K}_{j}$. We also introduce $\Theta_{i}$, a $d\times r_{i}$ matrix whose columns form an orthonormal basis for $\widetilde{Q_{i}}$, where $r_{i}$ is the dimension of $\widetilde{Q_{i}}$. With these definitions in place, we can formulate the SDP for determining the optimal probability of unambiguous discrimination of the ensemble $\mathcal{E}$ as follows
Primal problem
maximize: $\sum_{i=1}^{N}q_{i}\operatorname{Tr}(\sigma_{i}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger})$
subject to: $\sum_{i=1}^{N}\Theta_{i}\Delta_{i}\Theta_{i}^{\dagger}\preceq\mathbb{1},$ (8a)
$\Delta_{i}\succeq 0,\quad\forall i\in[N]$
Dual problem
minimize: $\operatorname{Tr}(Z)$
subject to: $\Theta_{i}^{\dagger}\left(Z-q_{i}\sigma_{i}\right)\Theta_{i}\succeq 0\quad\forall i\in[N],$
$Z\succeq 0,$ (8b)
where $\Delta_{i}$ is an $r_{i}\times r_{i}$ matrix for each $i\in[N]$. The optimization variables in this formulation are the $N$ matrices denoted by $\Delta_{i}$. Our proof strategy will be to use Slater’s theorem which states that if the primal problem is convex and strict feasibility holds, then the duality gap is zero and the primal and dual optimal values are equal boyd2004convex . We’ll first assume primal and dual optimal variables for the ensemble SDPs. Then we will use these variables to construct primal and dual feasible solutions for the sequence SDP. Finally we will show that these variables make the primal and dual value equal. Applying Slater’s theorem to the sequence SDP, we will conclude that these must be the optimal solutions.
Proof.
Consider the ensemble
$\mathcal{E}^{i}=\{(\eta^{i}_{1},\rho^{i}_{1}),\ldots,(\eta^{i}_{\ell_{i}},\rho^{i}_{\ell_{i}})\}$ (9)
where $\rho_{j}^{i}$ are $d_{i}\times d_{i}$ density operators and $j\in[\ell_{i}]$. Let $S^{i}_{j}=S(\mathcal{E}^{i})\setminus\{\rho^{i}_{j}\}$ for some $j\in[\ell_{i}]$ and $\widetilde{S^{i}_{j}}$ be the intersection of the kernels of all the density matrices of $\mathcal{E}^{i}$, except for $\rho^{i}_{j}$. That is, $\widetilde{S^{i}_{j}}=\cap_{t=1,t\neq j}^{\ell_{i}}\mathcal{K}_{t}^{i}$ where $\mathcal{K}_{t}^{i}$ is the kernel of $\rho_{t}^{i}$. Let $\Theta_{j}^{i}$ be a $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $\widetilde{S^{i}_{j}}$ (which is of dimension $r_{j}^{i}$). The optimal probability of unambiguous discrimination of the elements of $\mathcal{E}^{i}$ is denoted by $p(\mathcal{E}^{i})$ and let $\Delta_{j}^{i}$ be the $r_{j}^{i}\times r_{j}^{i}$ matrices that achieve this optimum. In addition to being positive, these matrices satisfy $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger}\preceq\mathbb{1}$. (See SM for the SDPs of the individual ensembles and the sequence.) We denote the optimal dual variable for ensemble $\mathcal{E}^{i}$ by $Z_{i}$, which is also a positive matrix and satisfies ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0$. The optimal primal and dual variables satisfy
$\sum_{j=1}^{\ell_{i}}\eta_{j}^{i}\operatorname{Tr}\left(\rho_{j}^{i}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger}\right)=\operatorname{Tr}(Z_{i})=p(\mathcal{E}^{i}).$ (10)
Now, we turn to the sequence ensemble given by equation (2) and consider the SDP of its unambiguous discrimination. Note that there are $\ell$ number of sequence states in this ensemble, and a state is labeled by the index $(x_{1},\ldots,x_{k})$, which is a $k$-tuple of indices where $x_{i}\in[\ell_{i}]$ for all $i\in[k]$. Let us denote by $\widetilde{S}(x_{1},\ldots,x_{k})$ the intersection of the kernels of all the states of $\mathcal{E}_{k}$ except for $\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}$. That is,
$\widetilde{S}(x_{1},\ldots,x_{k})=\bigcap_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\in[\ell_{1}]\times\cdots\times[\ell_{k}]\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\mathcal{K}(y_{1},\ldots,y_{k})$
where $\mathcal{K}(y_{1},\ldots,y_{k})$ is the kernel of $\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}}$. This subspace has a decomposition (see Theorem 9 in SM) as follows
$\widetilde{S}(x_{1},\ldots,x_{k})=\widetilde{S_{x_{1}}^{1}}\otimes\cdots\otimes\widetilde{S_{x_{k}}^{k}}$
$=(\cap_{j=1,j\neq x_{1}}^{\ell_{1}}\mathcal{K}_{j}^{1})\otimes\cdots\otimes(\cap_{j=1,j\neq x_{k}}^{\ell_{k}}\mathcal{K}_{j}^{k}).$ (11)
We first give a prescription to construct matrices $\Theta(x_{1},...,x_{k})$ whose columns form an orthonormal basis of $\widetilde{S}(x_{1},\ldots,x_{k})$. From equation (Proof.), it can be seen that these matrices are of size $\prod_{i=1}^{k}d_{i}\times\prod_{i=1}^{k}r^{i}_{x_{i}}(x_{i}\in[\ell_{i}])$ and a simple procedure to construct them is to take $\Theta(x_{1},...,x_{k})=\Theta_{x_{1}}^{1}\otimes\cdots\otimes\Theta_{x_{k}}^{k}$. For a feasible primal variable $\Delta(x_{1},...,x_{k})$, we take the tensor product of the optimal primal variables, $\Delta(x_{1},...,x_{k})=\Delta_{x_{1}}^{1}\otimes\cdots\otimes\Delta_{x_{k}}^{k}$. This operator is positive since it is the tensor product of positive operators. Also, the defined operators satisfy (using Lemma 3),
$\sum_{(x_{1},...,x_{k})}\Theta(x_{1},...,x_{k})\Delta(x_{1},...,x_{k})\Theta(x_{1},...,x_{k})^{\dagger}$ (12)
$=$ $\sum_{(x_{1},...,x_{k})}(\Theta_{x_{1}}^{1}\Delta_{x_{1}}^{1}{\Theta_{x_{1}}^{1}}^{\dagger})\otimes\cdots\otimes(\Theta_{x_{k}}^{k}\Delta_{x_{k}}^{k}{\Theta_{x_{k}}^{k}}^{\dagger})$
$\preceq$  $\;\mathbb{1}.$
Now consider the dual variable $Z=Z_{1}\otimes\cdots\otimes Z_{k}$, which is positive due to the positivity of the operators $Z_{i}$ for all $i\in[k]$. We also know that these variables satisfy ${\Theta_{x_{i}}^{i}}^{\dagger}Z_{i}\Theta_{x_{i}}^{i}\succeq{\Theta_{x_{i}}^{i}}^{\dagger}(\eta_{x_{i}}^{i}\rho_{x_{i}}^{i})\Theta_{x_{i}}^{i}$ for all $x_{i}\in[\ell_{i}]$ and $i\in[k]$. By taking tensor products for the indices $(x_{1},...,x_{k})$ we get (using Lemma 3)
$\Theta(x_{1},...,x_{k})^{\dagger}(Z-\eta_{x_{1}}\cdots\eta_{x_{k}}\rho_{x_{1}}\otimes\cdots\otimes\rho_{x_{k}})\Theta(x_{1},...,x_{k})\succeq 0.$ (13)
This shows that the primal and dual variables we introduced are feasible, and they both make the respective objective value equal to $\prod_{i}p(\mathcal{E}^{i})$ by the trace property of the tensor product. Therefore, $p(\mathcal{E}_{k})$ must be equal to $\prod_{i}p(\mathcal{E}^{i})$. ∎
Conclusions.—Sequence discrimination problems arise in various quantum information processing tasks, where the objective is to determine the state of a sequence of finite length. The main question here is whether an optimal measurement is local or collective, i.e., whether it suffices to measure the individual members of the sequence or not. In this paper, we showed that as long as the members of the sequences are drawn independently from the same ensemble or even from different ensembles, the optimal measurement in both minimum-error and unambiguous discrimination paradigms is local and comprises fixed measurements on the individual components. It follows that if we give up the condition of independent selection of the members of a sequence, the optimal measurement could be either local and fixed, local and adaptiveacin2005multiple , or collective (partially or wholly) calsamiglia2010local .
Our result also shows that independent sequences do not exhibit nonlocality in the sense some other unentangled systems do, as collective measurement is not necessary for optimal extraction of information from such sequences. However, in the context of quantum state exclusion, one may consider them nonlocal, for an entangled measurement is necessary, as demonstrated by the PBR result pusey2012reality and its generalizations crickmore2020unambiguous .
A significant part of our proof relied on solving the SDP formulation of the problem using the strong duality theorem boyd2004convex . While SDP has been successfully used to solve various problems in quantum information theory skrzypczyk2023semidefinite ; mironowicz2024semi , we believe our approach could help to solve problems in many-body unentangled systems.
Our result could be immediately applied to settings of quantum key distribution protocols Bennett_2014 ; bennett1992quantum or similar ones. In these protocols, Alice sends a sequence of quantum states - selected independently from a known ensemble - to Bob, who measures each state individually. However, if Bob can store the incoming states in quantum memory, he may think about performing a collective measurement on the entire sequence to extract more information. Our result rules out this possibility, as these sequences are i.i.d.
A natural direction for future research involves doing away with the independence assumption. Can there be sequences of dependent quantum states whose optimal discrimination is still achievable by fixed, local measurements? A systematic study involving the set of states, the nature of the joint probability distribution for the states of a sequence and the nature of optimal measurement to discriminate them will be interesting to investigate.
Acknowledgements.
The authors would like to thank Snehasish Roy Chowdhury and Jamie Sikora for fruitful discussions.
Appendix A Supplemental material
Appendix B Proof of Lemma 3
Proof.
Given that $A\succeq C$ and $B\succeq D$, we have that $A-C$ and $B-D$ are positive semidefinite. To show $A\otimes B\succeq C\otimes D$, consider the expression
$A\otimes B-C\otimes D$ $=(A\otimes B)-(A\otimes D)+(A\otimes D)-(C\otimes D)$ (14)
$=(A\otimes(B-D))+((A-C)\otimes D).$
Since $A-C$ and $B-D$ are positive semidefinite, their tensor products with any positive semidefinite operator also yield positive semidefinite operators. Hence, $A\otimes(B-D)$ and $(A-C)\otimes D$ are positive semidefinite. The sum of two positive semidefinite operators remains positive semidefinite. Therefore, $A\otimes B-C\otimes D$ is positive semidefinite, thus $A\otimes B\succeq C\otimes D$. ∎
Appendix C Proof of Lemma 5
We will need the following two lemmas to prove our claim.
Lemma 7.
Let $k$ be a positive integer and $\mathcal{V}$ be a vector space. For each $i\in[k]$, let $\mathcal{V}_{i}\subseteq\mathcal{V}$ and $\mathcal{W}_{i}\subseteq\mathcal{V}_{i}$ be subspaces. If $a_{i}\in\mathcal{V}_{i}$ and $a_{i}\notin\mathcal{W}_{i}$, for each $i\in[k]$, then $a_{1}\otimes\cdots\otimes a_{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$.
Proof.
For each $i\in[k]$, let $\{u_{1}^{i},\ldots,u_{s_{i}}^{i}\}$ be a basis of $\mathcal{W}_{i}$, and extend it to a basis $\{u_{1}^{i},\ldots,u_{s_{i}}^{i},v_{1}^{i},\ldots,v_{t_{i}}^{i}\}$ of $\mathcal{V}_{i}$. First, express each $a_{i}$ in terms of the basis of $\mathcal{V}_{i}$:
$a_{i}=b_{i}+\sum_{j=1}^{t_{i}}\lambda_{j}^{i}v_{j}^{i},$ (15)
where $b_{i}\in\mathcal{W}_{i}$. Since $a_{i}\notin\mathcal{W}_{i}$, there exists at least one $j_{i}$ such that $\lambda_{j_{i}}^{i}\neq 0$. Consider the tensor product $a_{1}\otimes\cdots\otimes a_{k}$. Its expansion includes the term:
$(\lambda_{j_{1}}^{1}\cdots\lambda_{j_{k}}^{k})(v_{j_{1}}^{1}\otimes\cdots\otimes v_{j_{k}}^{k}).$ (16)
This term is non-zero because each $\lambda_{j_{i}}^{i}\neq 0$. Observe that $v_{j_{1}}^{1}\otimes\cdots\otimes v_{j_{k}}^{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$, as each $v_{j_{i}}^{i}\notin\mathcal{W}_{i}$. Finally, note that $a_{1}\otimes\cdots\otimes a_{k}\notin\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$, because its expansion contains a non-zero term that is not in $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$. ∎
Lemma 8.
Let $\mathcal{V}_{i}$ be vector spaces and $B_{i}=\{b_{1}^{i},\ldots,b_{\ell_{i}}^{i}\}\subseteq\mathcal{V}_{i}$ for $i\in[k]$. Let $\mathcal{W}_{i}=\text{span}(B_{i})$ be the subspaces spanned by $B_{i}$. Then
$\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}=\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}.$ (17)
Proof.
We will show that $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}=\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}$ by proving both inclusions. First, consider
$\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}\subseteq\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}.$ (18)
For all $i\in[k]$ and each $x_{i}\in[\ell_{i}]$, we have $b_{x_{i}}^{i}\in\mathcal{W}_{i}$. Therefore, $b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\in\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$. As $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$ is a subspace, it contains the span of these tensors.
Now, for the reverse inclusion
$\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}\subseteq\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}.$ (19)
Let $w_{1}\otimes\cdots\otimes w_{k}$ be an arbitrary pure tensor in $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$. For each $i\in[k]$, we can write $w_{i}=\sum_{j=1}^{\ell_{i}}\lambda_{j}^{i}b_{j}^{i}$ as $w_{i}\in\mathcal{W}_{i}=\text{span}(B_{i})$. Expanding the pure tensor
$w_{1}\otimes\cdots\otimes w_{k}=\sum_{j_{1}=1}^{\ell_{1}}\cdots\sum_{j_{k}=1}^{\ell_{k}}(\lambda_{j_{1}}^{1}\cdots\lambda_{j_{k}}^{k})(b_{j_{1}}^{1}\otimes\cdots\otimes b_{j_{k}}^{k})$ (20)
This expansion shows that $w_{1}\otimes\cdots\otimes w_{k}\in\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}$. As $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}$ is spanned by such pure tensors, the inclusion follows. Therefore, $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{k}=\text{span}\{b_{x_{1}}^{1}\otimes\cdots\otimes b_{x_{k}}^{k}\mid x_{i}\in[\ell_{i}],\ \forall i\in[k]\}$. ∎
Proof of Lemma 5.
We first prove that if $S(\mathcal{E}^{i})$ satisfies Lemma 4 for each $i$, then $S(\mathcal{E}_{k})$ also satisfies Lemma 4. Therefore, we assume that for each $i\in[k]$, $S(\mathcal{E}^{i})$ satisfies Lemma 4. For each $\rho_{j}^{i}\in S(\mathcal{E}^{i})$, let $A(i,j)=\{|\psi(i,j,t)\rangle\}_{t=1}^{m_{ij}}$ be the set of eigenvectors corresponding to nonzero eigenvalues of $\rho_{j}^{i}$, where $m_{ij}$ is the number of such eigenvectors. Thus, $\text{supp}(\rho_{j}^{i})=\text{span}(A(i,j))$. Define
$A^{i}$ =$=\{|\psi(i,j,t)\rangle:t\in[m_{ij}],\ j\in[\ell_{i}]\},$ (21)
$A^{i}_{y}$ =$=\{|\psi(i,j,t)\rangle:t\in[m_{ij}],\ j\in[\ell_{i}],\ j\neq y\}\text{ for }y\in[\ell_{i}].$
By our assumption, for each $i\in[k]$, $\text{span}(A^{i})\neq\text{span}(A^{i}_{y})$ for all $y\in[\ell_{i}]$. Thus, there exists $t_{i,y}\in[m_{iy}]$ such that $|\psi(i,y,t_{i,y})\rangle\notin\text{span}(A^{i}_{y})$. Suppose, for contradiction, that $S(\mathcal{E}_{k})$ does not satisfy Lemma 4. Then there exist $y_{i}\in[\ell_{i}]$ for all $i\in[k]$ such that
$\text{supp}(S(\mathcal{E}_{k}))=\text{supp}(S(\mathcal{E}_{k})\setminus\{\rho_{y_{1}}^{1}\otimes\cdots\otimes\rho_{y_{k}}^{k}\}).$ (22)
The eigenvectors corresponding to non-zero eigenvalues of $\rho_{y_{1}}^{1}\otimes\cdots\otimes\rho_{y_{k}}^{k}$ are
$\{|\psi(1,y_{1},s_{1,y_{1}})\rangle\otimes\cdots\otimes|\psi(k,y_{k},s_{k,y_{k}})\rangle:s_{i,y_{i}}\in[m_{i,y_{i}}]\}.$ (23)
Thus,
$\text{supp}(S(\mathcal{E}_{k})\setminus\{\rho_{y_{1}}^{1}\otimes\cdots\otimes\rho_{y_{k}}^{k}\})$ (24)
$=\text{span}(\{|\psi(1,x_{1},s_{1,x_{1}})\rangle\otimes\cdots\otimes|\psi(k,x_{k},s_{k,x_{k}})\rangle:s_{i,x_{i}}\in[m_{i,x_{i}}],\ x_{i}\neq y_{i}\;\forall i\in[k]\})$
$=\bigotimes_{i=1}^{k}\text{span}(\{|\psi(i,x_{i},s_{i,x_{i}})\rangle\}),\;s_{i,x_{i}}\in[m_{i,x_{i}}],\ x_{i}\neq y_{i}\;\forall i\in[k]$
$=\text{span}(A^{1}_{y_{1}})\otimes\cdots\otimes\text{span}(A^{k}_{y_{k}}),$
where we have used Lemma 7 in the second equality. This equals $\text{supp}(S(\mathcal{E}_{k}))$ by our assumption. However by Lemma 8 we see that,
$|\psi(1,y_{1},t_{1,y_{1}})\rangle\otimes\cdots\otimes|\psi(k,y_{k},t_{k,y_{k}})\rangle\notin\text{span}(A^{1}_{y_{1}})\otimes\cdots\otimes\text{span}(A^{k}_{y_{k}}),$ (25)
leading to a contradiction.
We now prove the converse by showing that if for some $i\in[k],~{}S(\mathcal{E}^{i})$ does not satisfy Lemma 4 then $S(\mathcal{E}_{k})$ also fail to satisfy Lemma 4. Therefore, assume that for some $i\in[k]$, the set $S(\mathcal{E}^{i})$ does not satisfy Lemma 4. Then there exists a $y\in[\ell_{i}]$ such that $\text{span}(A^{i})=\text{span}(A^{i}_{y})$. This implies $|\psi(i,y,s_{i,y})\rangle\in\text{span}(A^{i}_{y})$ for all $s_{i,y}\in[m_{i,y}]$. Consider
$S(\mathcal{E}_{k})\setminus\{\rho_{1}^{1}\otimes\cdots\otimes\rho_{y}^{i}\otimes\cdots\otimes\rho_{1}^{k}\}.$ (26)
The support of the removed state is
$\text{span}\{|\psi(1,1,s_{1,1})\rangle\otimes\cdots\otimes|\psi(i,y,s_{i,y})\rangle\otimes\cdots\otimes|\psi(k,1,s_{k,1})\rangle:s_{a,b}\in[m_{a,b}]\}.$ (27)
Each element
$|\psi(1,1,s_{1,1})\rangle\otimes\cdots\otimes|\psi(i,y,s_{i,y})\rangle\otimes\cdots\otimes|\psi(k,1,s_{k,1})\rangle$ (28)
is in $\text{span}(A^{1}_{x_{1}})\otimes\cdots\otimes\text{span}(A^{i}_{y})\otimes\cdots\otimes\text{span}(A^{k}_{x_{k}})$ where $x_{j}\neq 1$ for all $j\in[k]$ except for $j=i$, for which $x_{j}=y$. However,
$\text{span}(A^{1}_{x_{1}})\otimes\cdots\otimes\text{span}(A^{i}_{y})\otimes\cdots\otimes\text{span}(A^{k}_{x_{k}})\subset\text{supp}(S(\mathcal{E}_{k})\setminus\{\rho_{1}^{1}\otimes\cdots\otimes\rho_{y}^{i}\otimes\cdots\otimes\rho_{1}^{k}\}).$ (29)
This shows that
$\text{supp}(S(\mathcal{E}_{k})\setminus\{\rho_{1}^{1}\otimes\cdots\otimes\rho_{y}^{i}\otimes\cdots\otimes\rho_{1}^{k}\})=\text{supp}(S(\mathcal{E}_{k})),$ (30)
proving that $S(\mathcal{E}_{k})$ fails to satisfy Lemma 4. ∎
Appendix D Proof of Theorem 9
Theorem 9.
Let $k\in\mathbb{N}$, and for each $i\in[k]$, let $S(\mathcal{E}^{i})=\{\rho_{1}^{i},\ldots,\rho_{\ell_{i}}^{i}\}$ be a set of $\ell_{i}$ quantum states which are density operators on $\mathbb{C}^{d_{i}}$ for some integer $d_{i}$. Define $S(\mathcal{E}_{k})=\{\rho_{x_{1}}^{1}\otimes\cdots\otimes\rho_{x_{k}}^{k}:x_{i}\in[\ell_{i}],\ \forall i\in[k]\}$ as the set of all $k$-length tensor product sequences where the $i$-th component comes from $S(\mathcal{E}^{i})$.
For each $i\in[k]$ and $j\in[\ell_{i}]$, let $\widetilde{S^{i}_{j}}$ be defined as
$\widetilde{S^{i}_{j}}=\bigcap_{m\in[\ell_{i}],m\neq j}\operatorname{ker}(\rho^{i}_{m}).$ (31)
For $(x_{1},\ldots,x_{k})\in[\ell_{1}]\times\cdots\times[\ell_{k}]$, let $\widetilde{S}(x_{1},\ldots,x_{k})$ be defined as
$\widetilde{S}(x_{1},\ldots,x_{k})=\bigcap_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\in[\ell_{1}]\times\cdots\times[\ell_{k}]\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{ker}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}}).$ (32)
Then, for all $(x_{1},\ldots,x_{k})\in[\ell_{1}]\times\cdots\times[\ell_{k}]$, it holds that
$\widetilde{S}(x_{1},\ldots,x_{k})=\widetilde{S^{1}_{x_{1}}}\otimes\cdots\otimes\widetilde{S^{k}_{x_{k}}}.$ (33)
Proof.
We expand both sides of Equation (33) in terms of the kernels. The right side can be written as
$\widetilde{S^{1}_{x_{1}}}\otimes\cdots\otimes\widetilde{S^{k}_{x_{k}}}$ $=\left(\bigcap_{j=1,j\neq x_{1}}^{\ell_{1}}\operatorname{ker}(\rho^{1}_{j})\right)\otimes\cdots\otimes\left(\bigcap_{j=1,j\neq x_{k}}^{\ell_{k}}\operatorname{ker}(\rho^{k}_{j})\right)$ (34)
$=\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right).$
The second equality follows from the fact that if $\{\mathcal{V}^{i}_{j}\}_{j=1}^{\ell_{i}}$ are subspaces of $\mathcal{V}^{i}$ for $i\in[k]$, then it holds that
$\left(\bigcap_{j=1}^{\ell_{1}}\mathcal{V}^{1}_{j}\right)\otimes\cdots\otimes\left(\bigcap_{j=1}^{\ell_{k}}\mathcal{V}^{k}_{j}\right)=\bigcap_{(j_{1},\ldots,j_{k})\in[\ell_{1}]\times\cdots\times[\ell_{k}]}\mathcal{V}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{V}^{k}_{j_{k}}.$ (35)
The left side of Equation (33) can be expanded as follows
$\widetilde{S}(x_{1},\ldots,x_{k})$ $=\bigcap_{\begin{subarray}{c}(y_{1},\ldots,y_{k})=(1,\ldots,1)\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}^{(\ell_{1},\ldots,\ell_{k})}\operatorname{ker}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})$ (36)
$=\bigcap_{\begin{subarray}{c}(y_{1},\ldots,y_{k})=(1,\ldots,1)\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}^{(\ell_{1},\ldots,\ell_{k})}\left(\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\mathbb{C}^{d_{2}}\otimes\cdots\otimes\mathbb{C}^{d_{k}}+\cdots+\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{k-1}}\otimes\operatorname{ker}(\rho^{k}_{y_{k}})\right).$
The second equality follows from the fact that if $\varphi_{i}$ is a linear operator on vector space $\mathcal{V}_{i}$ for $i\in[k]$ and its kernel is denoted by $\operatorname{ker}(\varphi_{i})$, then the following relation holds:
$\operatorname{ker}(\varphi_{1}\otimes\cdots\otimes\varphi_{k})=\sum_{i=1}^{k}\mathcal{V}_{1}\otimes\cdots\otimes\operatorname{ker}(\varphi_{i})\otimes\cdots\otimes\mathcal{V}_{k}.$ (37)
Now for each $k$-tuple $(y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})$, consider $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\mathbb{C}^{d_{2}}\otimes\cdots\otimes\mathbb{C}^{d_{k}}+\cdots+\mathbb{C}^{d_{1}}\otimes\cdots\otimes\mathbb{C}^{d_{k-1}}\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$. This sum contains $\operatorname{ker}(\rho^{1}_{y_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{y_{k}})$ and thus contains $\bigcap\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)$ where $j_{i}\neq x_{i}$ for all $i\in[k]$. Therefore, we conclude:
$\widetilde{S}(x_{1},\ldots,x_{k})\supseteq\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right).$ (38)
For the converse, we will show that the following relation holds
$\left(\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)\right)^{\perp}\subseteq\left(\widetilde{S}(x_{1},\ldots,x_{k})\right)^{\perp},$ (39)
where for a subspace $\mathcal{V}$, $\mathcal{V}^{\perp}$ denotes its orthogonal complement. First, we express the right side of the above relation as follows:
$\left(\widetilde{S}(x_{1},\ldots,x_{k})\right)^{\perp}$ $=\left(\bigcap_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{ker}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})\right)^{\perp}$ (40)
$=\sum_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{ker}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})^{\perp}$
$=\sum_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{Im}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})$
where $\operatorname{Im}(\rho)$ denotes the image of an operator $\rho$. In the second equality, we have used the fact that the orthogonal complement of the intersection of subspaces is equal to the sum of the orthogonal complement of individual subspaces: $(\mathcal{V}_{1}\cap\cdots\cap\mathcal{V}_{n})^{\perp}=\mathcal{V}_{1}^{\perp}+\cdots+\mathcal{V}_{n}^{\perp}$ where $\mathcal{V}_{i}$’s are subspaces. This identity also allows us to write the left side of Equation (39) as a sum of subspaces
$\left(\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)\right)^{\perp}=\sum_{\begin{subarray}{c}(j_{1},\ldots,j_{k})\\ j_{i}\neq x_{i},\forall i\end{subarray}}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)^{\perp}.$ (41)
This can be further expanded by noting that for subspaces $\mathcal{V}_{1},\ldots,\mathcal{V}_{n}$, we have
$(\mathcal{V}_{1}\otimes\cdots\otimes\mathcal{V}_{n})^{\perp}=$ $~{}\mathcal{V}_{1}^{\perp}\otimes\cdots\otimes\mathcal{V}_{n}+\cdots+\mathcal{V}_{1}\otimes\cdots\otimes\mathcal{V}_{n}^{\perp}$ (42)
$+$ $~{}\mathcal{V}_{1}^{\perp}\otimes\mathcal{V}_{2}^{\perp}\otimes\cdots\otimes\mathcal{V}_{n}+\cdots+\mathcal{V}_{1}\otimes\cdots\mathcal{V}_{n-1}^{\perp}\otimes\mathcal{V}_{n}^{\perp}$
[…]
$+$ $~{}\mathcal{V}_{1}^{\perp}\otimes\cdots\otimes\mathcal{V}_{n}^{\perp}.$
Each term in the above sum is of the form $\mathcal{W}_{1}\otimes\cdots\otimes\mathcal{W}_{n}$ where $\mathcal{W}_{i}$ is either $\mathcal{V}_{i}$ or $\mathcal{V}_{i}^{\perp}$, and there is at least one index $a\in[n]$ for which $\mathcal{W}_{a}=\mathcal{V}_{a}^{\perp}$. Therefore, each term on the right side of Equation (41) is a sum of terms like $(\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}})$. That is,
$(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}}))^{\perp}=\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ (43)
where $\mathcal{L}^{i}_{j_{i}}=\operatorname{ker}(\rho^{i}_{j_{i}})$ or $\mathcal{L}^{i}_{j_{i}}=(\operatorname{ker}(\rho^{i}_{j_{i}}))^{\perp}$ and there is at least one index $a\in[k]$ for which $\mathcal{L}^{a}_{j_{a}}=(\operatorname{ker}(\rho^{a}_{j_{a}}))^{\perp}$.
Let us now consider a pure tensor $u_{1}\otimes\cdots\otimes u_{k}\in\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$, where $u_{a}\in(\operatorname{ker}(\rho^{a}_{j_{a}}))^{\perp}=\operatorname{Im}(\rho^{a}_{j_{a}})$ for some $a\in[k]$ and $j_{a}\neq x_{a}$. This means that $u_{a}=\rho^{a}_{j_{a}}(v^{a}_{j_{a}})$ for some $v^{a}_{j_{a}}\in\mathbb{C}^{d_{a}}$. For some other index $b$, where $\mathcal{L}^{b}_{j_{b}}=\operatorname{ker}(\rho^{b}_{j_{b}})$, $u_{b}$ can be written as $\sum_{i_{b}=1}^{\ell_{b}}\rho^{b}_{i_{b}}(v^{b}_{i_{b}})$ where $v^{b}_{i_{b}}\in\mathbb{C}^{d_{b}}$ for all $i_{b}$. This follows from the assumption that $\text{supp}(A^{i})$ (the span of all the eigenvectors of the states of $S^{i}$ that correspond to nonzero eigenvalues) spans $\mathbb{C}^{d_{i}}$ for all $i$. Therefore, this pure tensor can be written as
$~{}u_{1}\otimes\cdots\otimes u_{k}$ (44)
$=$ $\sum_{i=1}^{\ell_{1}}\rho^{1}_{i}(v^{1}_{i})\otimes\cdots\otimes\rho^{a}_{j_{a}}(v^{a}_{j_{a}})\otimes\cdots\otimes\sum_{i=1}^{\ell_{k}}\rho^{k}_{i}(v^{k}_{i})$
where the sum appears in those places whose corresponding $\mathcal{L}$ equals the kernel. This can be written as
$\sum\;\rho^{1}_{j_{1}}(v^{1}_{j_{1}})\otimes\cdots\otimes\rho^{a}_{j_{a}}(v^{a}_{j_{a}})\otimes\cdots\otimes\rho^{k}_{j_{k}}(v^{k}_{j_{k}})$ (45)
$=$ $\sum\;(\rho^{1}_{j_{1}}\otimes\cdots\otimes\rho^{a}_{j_{a}}\otimes\cdots\otimes\rho^{k}_{j_{k}})(v^{1}_{j_{1}}\otimes\cdots\otimes v^{a}_{j_{a}}\otimes\cdots v^{k}_{j_{k}})$
where the sum is over those indices whose corresponding $\mathcal{L}$ is equal to the kernel.
Therefore, the pure tensor, which was assumed to be an element of $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ is shown to belong to
$\sum_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{Im}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})$ (46)
since there is at least one index $a\in[k]$ that satisfies $j_{a}\neq x_{a}$.
Now observe that $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$ is spanned by pure tensors and since
$\left(\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)\right)^{\perp}$ (47)
is sum of spaces of the form $\mathcal{L}^{1}_{j_{1}}\otimes\cdots\otimes\mathcal{L}^{k}_{j_{k}}$, it is spanned by pure tensors as well. The span of these pure tensors form a subspace in
$\sum_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{Im}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})$ (48)
and any element of
$\left(\bigcap_{j_{i}\neq x_{i},\forall i}\left(\operatorname{ker}(\rho^{1}_{j_{1}})\otimes\cdots\otimes\operatorname{ker}(\rho^{k}_{j_{k}})\right)\right)^{\perp}$ (49)
also belongs to this subspace and hence to
$\sum_{\begin{subarray}{c}(y_{1},\ldots,y_{k})\\ (y_{1},\ldots,y_{k})\neq(x_{1},\ldots,x_{k})\end{subarray}}\operatorname{Im}(\rho^{1}_{y_{1}}\otimes\cdots\otimes\rho^{k}_{y_{k}})=\left(\widetilde{S}(x_{1},\ldots,x_{k})\right)^{\perp}.$ (50)
Therefore, the converse is established. ∎
Appendix E Semidefinite program of unambiguous discrimination of sequences
The SDPs of the ensembles and the sequences follow straightforwardly from the SDP given in Equation (8) in the main text. However, for the reader’s convenience, we write them here explicitly.
The SDP for the optimal probability of unambiguous discrimination of the states of ensemble $\mathcal{E}^{i}$ is the following
Primal problem maximize: $\sum_{j=1}^{\ell_{i}}\eta_{j}^{i}\operatorname{Tr}(\rho_{j}^{i}\Theta_{j}^{i}\Delta_{j}^{i}{\Theta_{j}^{i}}^{\dagger})$ subject to: $\sum_{j=1}^{\ell_{i}}\Theta_{j}^{i}\Delta_{j}^{i}\Theta_{j}^{i\dagger}\preceq\mathbb{1},$ $\Delta_{j}^{i}\succeq 0,\quad\forall j\in[\ell_{i}]$              Dual problem minimize: $\operatorname{Tr}(Z_{i})$ subject to: ${\Theta_{j}^{i}}^{\dagger}\left(Z_{i}-\eta_{j}^{i}\rho_{j}^{i}\right)\Theta_{j}^{i}\succeq 0\quad\forall j\in[\ell_{i}],$ $Z_{i}\succeq 0.$
Here $\Theta_{j}^{i}$ is an $d_{i}\times r_{j}^{i}$ matrix whose columns form an arbitrary orthonormal basis for $S_{j}^{i}$ (of dimension $r_{j}^{i}$).
The ensemble $\mathcal{E}_{k}$ consists of length $k$ sequences of quantum states that are chosen independently,
$\mathcal{E}_{k}=\{(\eta^{1}_{x_{1}}\cdots\eta^{k}_{x_{k}},\rho^{1}_{x_{1}}\otimes\dots\otimes\rho^{k}_{x_{k}}):~{}x_{i}\in[\ell_{i}]~{}\forall i\in[k]\}.$
The SDP for its optimal unambiguous discrimination is the following
Primal problem
maximize: $\sum_{x_{1}=1}^{\ell_{1}}\cdots\sum_{x_{k}=1}^{\ell_{k}}\eta_{x_{1}}^{1}\cdots\eta_{x_{k}}^{k}\operatorname{Tr}(\rho^{1}_{x_{1}}\otimes\cdots\otimes\rho^{k}_{x_{k}}\Theta(x_{1},...,x_{k})\Delta(x_{1},...,x_{k}){\Theta(x_{1},...,x_{k})}^{\dagger})$
subject to: $\sum_{x_{1}=1}^{\ell_{1}}\cdots\sum_{x_{k}=1}^{\ell_{k}}\Theta(x_{1},...,x_{k})\Delta(x_{1},...,x_{k}){\Theta(x_{1},...,x_{k})}^{\dagger}\preceq\mathbb{1},$ (52)
$\Delta(x_{1},...,x_{k})\succeq 0,\quad\forall x_{i}\in[\ell_{i}]~{}\forall i\in[k].$
Dual problem
minimize: $\operatorname{Tr}(Z)$
subject to: $\Theta(x_{1},...,x_{k})^{\dagger}(Z-\eta_{x_{1}}\cdots\eta_{x_{k}}\rho_{x_{1}}\otimes\cdots\otimes\rho_{x_{k}})\Theta(x_{1},...,x_{k})\succeq 0~{}\forall x_{i}\in[\ell_{i}]~{}\forall i\in[k],$ (53)
$Z\succeq 0.$
Here the matrices $\Theta(x_{1},...,x_{k})$ and $\Delta(x_{1},...,x_{k})$ are $\ell$ in number, one for each $k$-tuple $(x_{1},...,x_{k})$.